An initially motionless test car is accelerated uniformly to 1.10×102 km/h in 8.58 s before striking a simulated deer. The car is in contact with the faux fawn for 0.755 s, after which the car is measured to be traveling at 82.0 km/h. What is the magnitude of the acceleration of the car before the collision?

acceleration before collision:

What is the magnitude of the average acceleration of the car during the collision?

average acceleration during collision:

What is the magnitude of the average acceleration of the car during the entire test, from when the car first begins moving until the collision is over?

average acceleration during entire test:
Answer #1

acceleration before collision,

vf = 110 km/h = 30.55 m/s

vi = 0 m/s

a = change in velocity / time interval

a = vf - vi / t

a = 30.555 - 0 / 8.58

a = 3.56 m/s2

-----------------------------------------------------------------------------------

acceleration during collision:

v = 82 km/h = 22.77 m/s

a = 22.77 - 30.555 / 0.755

a = 10.3 m/s2

---------------------------------------------------------------------

average acceleration of the car during the entire test

a = 22.77 - 0 / 0.755 + 8.58

a = 2.439 m/s2